The table below shows data for the production of Avocados for an individual firm operating in a perfectly competitive market. Suppose that the price of Avocados is $6.

Number of workers Number of Avocados
0 0
10 90
20 162
30 216
40 252
50 270

Given this data, what is the marginal product of labor (MPL) when the quantity of workers goes from 0 to 10?
Answer #1

Marginal product of labor is the change in output that results from employing one additional unit of labor.

MPL = dQ / dL [ where dQ is the change in the number of avocados and dL is the change in number of workers]

If quantity of workers goes from 0 to 10, number of avocados Increases from 0 to 90.

Therefore, MPL = (90 - 0)/(10 - 0) = 90/10 = 9